OpenStudy (anonymous):

Can someone help me proof that (integral of sen^3 x dx from 0 to pi/4) is minus or equal to (integral of sen^2 x dx from 0 to pi/4) without doing the integrals?

myininaya (myininaya):

sen is a new function?

myininaya (myininaya):

what is its meaning?

OpenStudy (anonymous):

maybe you mean sine?

OpenStudy (anonymous):

sorry, sen is sine in portuguese

myininaya (myininaya):

lol

OpenStudy (anonymous):

sure. because \[\sin(x)\leq 1\] so \[\sin^3(x)\leq \sin^2(x)\]

myininaya (myininaya):

\[\int\limits_{0}^{\frac{\pi}{4}}\sin(x)(1-\cos^2(x)) dx\]

myininaya (myininaya):

oh without doing integrals

OpenStudy (anonymous):

inequality follows right?

OpenStudy (anonymous):

no one can do those integrals. they are too hard

OpenStudy (anonymous):

satellite73: dont understand "inequality follows right?"

OpenStudy (anonymous):

this is easy i think

OpenStudy (anonymous):

sine from 0 to pi/4 will always be between 0 and 1

OpenStudy (valpey):

It isn't necessary to do integrals. You just need to show that the cube of any number between 0 and 1 is less than or equal to the square of that same number.

OpenStudy (anonymous):

so as you multiply it by itself it gets smaller and smaller

OpenStudy (anonymous):

It is a consecuence of sin²(x)= 0.5(1-cos(2x))

OpenStudy (anonymous):

am i rite?

OpenStudy (valpey):

It follows from if f(x) <= g(x) for all x between a and b then the integral from a to b of f(x) is less than g(x). In this case f(x) is sin^3(x) and g(x) is sin^2(x).

OpenStudy (anonymous):

i mean inequality follows because if you have a positive number less that one, say x. then \[x^2<x\] for sure

OpenStudy (anonymous):

yea squeeze theorem ftw?

OpenStudy (anonymous):

Thanks so much for all of you that helped me with that one!

OpenStudy (anonymous):

point is that on the interval \[[0,\frac{\pi}{4}]\] \[\sin^3(x)\leq \sin^2(x)\] so \[\int_0^{\frac{\pi}{4}} \sin^3(x)dx \leq \int_0^{\frac{\pi}{4}} \sin^2(x)dx\]